# Synergistic eigenanalysis of covariance and Hessian matrices for enhanced binary classification

### v1

@@ 5 Discussion
Our work provides a compelling theoretical insight and a powerful, practical method, demonstrating the strength of simplicity in achieving remarkable results. Firstly, we provide a profound theoretical insight, revealing a subtle yet powerful relationship between covariance and Hessian matrices. Our formal proof seamlessly links covariance eigenanalysis with the first LDA criterion while Hessian eigenanalysis with the second one. This unification under LDA criteria offers a fresh and intuitive perspective on their interplay. Secondly, capitalizing on this theoretical elegance and simplicity, we introduce a novel method that consistently outperforms established techniques across diverse datasets. The unexpected efficacy of our method, rooted in the straightforward relationship between covariance, Hessian, and LDA, showcases the effectiveness of simplicity in addressing complex challenges.
@@
The proposed method outperforms PCA and the Hessian method by comprehensively addressing both LDA criteria—maximizing between-class mean distance and minimizing within-class variances. Unlike PCA, which predominantly focuses on the former and lacks the guidance of class labels, our supervised approach considers both aspects. Despite the computational efficiency associated with unsupervised dimension reduction methods (shen2014generalized, ), our approach demonstrates the added value of incorporating class labels. This key insight also underlies our outperformance of KPCA, which, despite operating on non-linearities, remains essentially unsupervised in nature. While the Hessian method concentrates on minimizing within-class variances, our method optimally combines the strengths of PCA and the Hessian, effectively identifying feature space directions that enhance both between-class separation and within-class compactness.
@@
While the proposed method employs the LDA criteria, it surpasses LDA itself in all cases. Figure 3 visually demonstrates the advantages of the proposed method over LDA. LDA is limited to a one-dimensional projection for binary classification problems ye2004two , where it seeks to identify a single direction that simultaneously satisfies the two criteria. Conversely, the proposed method splits the task of meeting the criteria into two directions. The utilization of higher dimensionality in the proposed method increases the likelihood of discovering class separability, aligning with Cover’s theorem (cover1965geometrical, ). KDA, operating on a non-linear mode, outperforms LDA in all cases, yet it remains fundamentally confined to one dimension, restricting its effectiveness in capturing intricate class-specific patterns compared to our proposed method.
Figure 3: Projection and classification results of UMAP, LDA, and the proposed method with SVM on the Wisconsin breast cancer test data. This figure showcases the test data projected by three distinct data projection methods, each separated by its corresponding SVM hyperplane trained using the WBCD training data. This figure highlights the effectiveness of linear SVM in facilitating class separability visualization and interpretability of the model.
@@
Our method outperforms LOL by leveraging non-linear modes of operation, providing a distinct advantage in capturing complex patterns beyond the linear capabilities of LOL. UMAP exhibits good class separability with the widest margin as shown in Figure 3. However, it is important to note that UMAP is not inherently a classification technique, and thus, it fails to generalize well to new data. Similar limitations apply to LLE, which, although effective in revealing local data structures, lacks the inherent capability for classification and generalization.
@@
Figure 3 also underscores the simplicity and interpretability of linear SVMs as basic linear classifiers in dealing with low-dimensional data. The figure shows clear visualizations of the SVM’s decision boundaries and separation achieved through the various dimensionality reduction methods. Notably, our proposed method combined with linear SVMs offers valuable insights into the decision-making process within the underlying DNN. Leveraging the transparency and interpretability provided by SVMs to address the inherent opacity of DNNs, our method bridges the gap between complex DNNs and comprehensible, highly accurate decision processes within a 2D space.
@@
The proposed method shows promising performance, but it has certain limitations. The applicability of our approach is based on the premise that a DNN model is already in place. Our method operates on top of the underlying DNN. Any shortcomings or biases in the DNN’s performance will naturally reflect in the results obtained from our approach. The limitation of our method lies in its dependence on the quality and accuracy of the underlying DNN.
@@
An important trajectory for future work involves investigating the extension of our methodology to accommodate various loss functions beyond binary cross-entropy. The mathematical derivation in our current work relies on the elegant relationship between binary cross-entropy loss and within-class variances. Exploring the adaptability of our method to different loss functions will contribute to a more comprehensive understanding of the method’s versatility, but requires careful scrutiny to establish analogous connections. Simultaneously, we recognize the need to extend our methodology from binary to multiclass classification. The binary classification focus in this work stems from foundational aspects guiding our formal proof, which is designed around binary assumptions to facilitate a streamlined and elegant derivation process. In particular, the use of binary cross-entropy as the loss function and the utilization of a linear SVM for evaluation inherently adhere to binary classification. Moving forward, careful exploration is needed to adapt our approach to multiclass scenarios to ensure its applicability and effectiveness across a broader range of classification tasks.
@@ 6 Conclusion
Our paper presents a multifaceted contribution to the realms of binary classification and dimensionality reduction. We offer a rigorous formal proof, showcasing how our novel method, grounded in the eigenanalysis of covariance and Hessian matrices, systematically enhances class separability, drawing inspiration from LDA criteria. This theoretical foundation finds strong empirical support through comprehensive experiments across an array of datasets. In these experiments, our approach consistently outperforms established techniques, demonstrating its robustness and applicability. Moreover, our approach shares a similar complexity profile with the traditional method, ensuring its practical utility in real-world scenarios. Additionally, through the integration of our method with linear SVMs, we improve the explainability of the intricate decision-making processes inherent to DNNs, addressing their typical lack of transparency and facilitating enhanced model interpretability.
@@ 7 Reproducibility statement
The detailed proofs for the theoretical foundations, emphasizing the maximization of between-class mean distance and minimization of within-class variance, are provided in Appendix A. The complete source code and datasets for our experiments are accessible through the following links:
@@ A.1 Proof of Theorem 1 - Maximizing Covariance for Maximizing Squared Between-Class Mean Distance
For any pair of data points $x_{i}$ and $x_{j}$, where $1\leq i\leq n$ and $n+1\leq j\leq 2n$, the difference between $x_{i}$ and $m_{1}$ is equal to the difference between $x_{j}$ and $m_{2}$, i.e., $x_{i}-m_{1}=x_{j}-m_{2}$. This relationship ensures that the shifted copies of $C_{1}$ and $C_{2}$ maintain the same relative distances from their respective means, preserving the identical distribution in both sets. Furthermore, the variances of $C_{1}$ and $C_{2}$, are equal, denoted as $s_{1}^{2}=s_{2}^{2}=s_{w}^{2}$, and the combined data from $C_{1}$ and $C_{2}$ has a variance of $s^{2}$.
@@
$s^{2}=\frac{1}{2n-1}\left(\sum_{i=1}^{n}(x_{i}-m)^{2}+\sum_{i=n+1}^{2n}(x_{i}-m)^{2}\right)$
@@
$s^{2}$ $=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)$
@@
$=\frac{1}{2}\left(s_{w}^{2}+s_{w}^{2}\right)+\frac{1}{4}d^{2}$
$=s_{w}^{2}+\frac{1}{4}d^{2}.$
Now, considering the data points representing the projected data onto an (Eigen)vector, we can utilize the Variance Ratio Preservation Theorem for Projection onto a Vector, which establishes the relationship between $s^{2}$ and $s_{w}^{2}$ as follows:
$s_{w}^{2}=\lambda\cdot s^{2}$
@@
$s^{2}$ $=\lambda\cdot s^{2}+\frac{1}{4}d^{2}.$
Let’s rearrange the equation by moving $2\lambda\cdot s^{2}$ to the left side:
$s^{2}-\lambda\cdot s^{2}$ $=\frac{1}{4}d^{2}.$
@@
$(1-\lambda)\cdot s^{2}$ $=\frac{1}{4}d^{2}.$
To solve for $s^{2}$, divide both sides by $(1-\lambda)$:
$s^{2}$ $=\frac{\frac{1}{4}d^{2}}{1-\lambda}=\frac{r^{2}}{1-\lambda}$
@@
We observe that the sign of $s^{2}$ and $d^{2}$ will be the same since the denominator $1-\lambda$ is always positive (as $10<\lambda<$ 1). Therefore, $s^{2}$ is linearly proportional to $d^{2}$.
Hence, maximizing the variance ($s^{2}$) will maximize the squared between-class mean distance ($d^{2}$) as desired.
@@ A.2 Proof of Theorem 2 - Maximizing Hessian for Minimizing Within-Class Variance
However, according to barshan2020relatif , we can approximate the Hessian using Fisher information:
$\mathrm{H}_{\theta}\approx\mathbb{E}_{p_{\theta}}\left[\left(\nabla_{\theta}\log p_{\theta}(c_{i}\mid\theta)\right)^{2}\right]$ (5)
@@
where $\mu$ and $\sigma$ are known constants, we can compute the Fisher information as follows:
$\mathbb{E}_{p_{\theta}}\left[\left(\nabla_{\theta}\log p_{\theta}(c_{i}\mid\theta)\right)^{2}\right]$ $=\mathbb{E}_{p_{\theta}}\left[\left(\nabla_{\theta}\log\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(\theta-\mu)^{2}}{2\sigma^{2}}\right)\right)^{2}\right]$
@@
From Eq(5), we derive the Hessian as follows:
$\mathrm{H}_{\theta}\approx\frac{1}{\sigma^{2}}$ (6)
@@
Combining the previous result with the Hessian calculation in Eq(6), we can conclude that:
$\mathrm{H}_{\theta}$ $\approx\frac{1}{\sigma_{\text{post}}^{2}}$

### v2

@@ 7 Discussion
Our work offers a valuable theoretical perspective and a practical method, emphasizing the utility of simplicity in achieving noteworthy outcomes. We present a theoretical framework that elucidates a significant relationship between covariance and Hessian matrices. This framework, supported by our formal proof, connects covariance eigenanalysis with the first LDA criterion (the concept of separation) and Hessian eigenanalysis with the second (the concept of compactness). While this theoretical clarity is compelling, the effectiveness of our method, demonstrated across diverse datasets, is contingent upon relatively ideal conditions as outlined in Section 3. These conditions include assumptions about data isotropy, specifically the dataset subsets corresponding to each class being approximately isotropic around their respective means, and the dominance of leading eigenvalues, which are crucial for the method’s success in practical applications.
The experiments conducted on the WBCD dataset illustrate these points. As shown in Figure 9, the absolute covariance matrices for each class in the WBCD dataset reveal a pattern that approximates the isotropic condition. The relatively uniform higher values along the diagonal and generally lower values in the off-diagonal elements indicate that the covariance matrices are nearly proportional to the identity matrix. As discussed in Section 3, the WBCD dataset being normalized (as with the other datasets used in this study) helps approximate isotropy, as does the dataset being a good dataset with low redundancy among attributes. This near-isotropy supports the method’s effectiveness in maximizing the between-class mean distance when projecting the WBCD data into the leading eigenvector of the covariance matrix, as suggested by Theorem 1.
Moreover, the eigenspectra displayed in Figure 10 highlight the dominance of the first eigenvalues in both the covariance and Hessian matrices. The eigenvalues are spread over several orders of magnitude on a logarithmic scale, often with approximately uniform spacing. Notably, the first eigenvalue for both matrices is situated on a different decade from the second and the subsequent eigenvalues indicating a significant separation in magnitude. This characteristic, especially pronounced in the Hessian matrix, is indicative of a ”sloppy” model [53, 54, 55, 56, 57, 58, 59], where a few stiff directions dominate. Similarly, in the covariance matrix, this pattern suggests a dominant underlying structure or trend in the data. This distribution ensures that the leading eigenvector captures a significant portion of the variance or curvature, crucial for the proposed method’s effectiveness in distinguishing between classes.
Figure 9: Absolute covariance matrices for WBCD dataset by class. The figure presents the absolute values of the covariance matrices for the WBCD dataset, calculated separately for each class. These matrices highlight the strength of linear relationships between features within each class. In these absolute covariance matrices, the approximate isotropic condition is evidenced by the relatively uniform higher values along the diagonal and the generally lower values in the off-diagonal elements. Specifically, for Class 0, the average diagonal value is 0.5655, and the average non-diagonal value is 0.1566; for Class 1, these values are 0.8503 and 0.2828, respectively. These characteristics, though not starkly isotropic, approximate the conditions where the covariance matrices are proportional to the identity matrix. This approximation helps ensure the applicability of Theorem 1 in higher dimensions.
Figure 10: Eigenspectra of Covariance and Hessian Matrices for the WBCD Dataset. The plots display the eigenspectra of the covariance (left) and Hessian (right) matrices on a logarithmic scale. In both spectra, it is visible that the eigenvalues are spread over several orders of magnitude, often with approximately uniform spacing between them. Remarkably, the first eigenvalue in both matrices is positioned an order of magnitude above the second and subsequent eigenvalues, indicating a substantial difference in scale. This distribution suggests that the leading eigenvector captures a substantial portion of the variance or curvature, which is crucial for the effectiveness of the proposed method.
In contrast, datasets that do not fit the ideal conditions of isotropy and dominant leading eigenvalues may present challenges for the proposed method. For instance, if the covariance matrix elements are random or lack a pattern approximating proportionality to the identity matrix, which can occur in unnormalized datasets or datasets with high redundancy among attributes, the method’s effectiveness may be compromised. Additionally, if the covariance and/or Hessian matrices exhibit relatively uniform eigenvalues, where the leading eigenvalues are not sufficiently dominant, the method may not perform as effectively.
@@
In our experiments, the proposed method outperforms PCA and the Hessian method by comprehensively addressing both LDA criteria—maximizing between-class mean distance and minimizing within-class variances. Unlike PCA, which predominantly focuses on the former and lacks the guidance of class labels, our supervised approach considers both aspects. Despite the computational efficiency associated with unsupervised dimension reduction methods [72], our approach demonstrates the added value of incorporating class labels. This key insight also underlies our outperformance of KPCA, which, despite operating on non-linearities, remains essentially unsupervised in nature. While the Hessian method concentrates on minimizing within-class variances, our method optimally combines the strengths of PCA and the Hessian, effectively identifying feature space directions that enhance both between-class separation and within-class compactness.
@@
While the proposed method employs the LDA criteria, it surpasses LDA itself in all cases. Figures 7 and 8 visually demonstrates the advantages of the proposed method over LDA. LDA is limited to a one-dimensional projection for binary classification problems [51], where it seeks to identify a single direction that simultaneously satisfies the two criteria. Conversely, the proposed method splits the task of meeting the criteria into two directions. The utilization of higher dimensionality in the proposed method increases the likelihood of discovering class separability, aligning with Cover’s theorem [73]. KDA, operating on a non-linear mode, outperforms LDA in all cases, yet it remains fundamentally confined to one dimension, restricting its effectiveness in capturing intricate class-specific patterns compared to our proposed method.
@@
Our method outperforms LOL by leveraging non-linear modes of operation, providing a distinct advantage in capturing complex patterns beyond the linear capabilities of LOL. UMAP exhibits good class separability with the widest margin as shown in Figures 7 and 8. However, it is important to note that UMAP is not inherently a classification technique, and thus, it fails to generalize well to new data in Figure 8 . Similar limitations apply to LLE, which, although effective in revealing local data structures, lacks the inherent capability for classification and generalization.
@@
Figures 7 and 8 also underscores the simplicity and interpretability of linear SVMs as basic linear classifiers in dealing with low-dimensional data. The figure shows clear visualizations of the SVM’s decision boundaries and separation achieved through the various dimensionality reduction methods. Notably, our proposed method combined with linear SVMs offers valuable insights into the decision-making process within the underlying DNN. Leveraging the transparency and interpretability provided by SVMs to address the inherent opacity of DNNs, our method bridges the gap between complex DNNs and comprehensible, highly accurate decision processes within a 2D space.
In another perspective, our proposed method highlights the effective use of DNNs for data projection and representation. In this approach, DNNs serve as a classifier whose decision-making process can be encapsulated in the Hessian matrix. By extracting the leading eigenvector from this matrix, we capture significant classification information, which is then used as one of the key projection directions in our proposed method. This process leverages the deep learning model’s ability to distill complex patterns into a compact and informative representation. This aligns with the findings of previous work, such as that by Hinton and Salakhutdinov [74], which highlights the potential of neural networks in enhancing data representation and separability.
@@
The proposed method shows promising performance, but it has certain limitations. The applicability of our approach is based on the premise that a DNN model is already in place. Our method operates on top of the underlying DNN, so any shortcomings or biases in the DNN’s performance will naturally reflect in the results obtained from our approach. The limitation of our method lies in its dependence on the quality and accuracy of the underlying DNN.
Another limitation of the proposed method arises from its reliance on relatively ideal dataset conditions for optimal performance. In practical applications, not all datasets will meet these criteria. For instance, datasets with high redundancy among attributes, lacking proper normalization, or involving models that are not sufficiently sloppy may not exhibit the desired isotropy or dominance of leading eigenvalues. In such cases, the method’s performance may degrade, potentially limiting its applicability to a broader range of datasets.
@@
An important trajectory for future work involves investigating the extension of our methodology to accommodate various loss functions beyond binary cross-entropy. The mathematical derivation in our current work relies on the elegant relationship between binary cross-entropy loss and within-class variances. Even though the binary cross-entropy loss function is widely recognized as the standard loss function for binary classification models [75], exploring the adaptability of our method to different loss functions will contribute to a more comprehensive understanding of the method’s versatility, but requires careful scrutiny to establish analogous connections.
Simultaneously, we recognize the need to extend our methodology from binary to multiclass classification. The binary classification focus in this work stems from foundational aspects guiding our formal proof, which is designed around binary assumptions to facilitate a streamlined and elegant derivation process. In particular, the use of binary cross-entropy as the loss function and the utilization of a linear SVM for evaluation inherently adhere to binary classification. Moving forward, careful exploration is needed to adapt our approach to multiclass scenarios to ensure its applicability and effectiveness across a broader range of classification tasks.
It is also important to validate our method on more complex and modern datasets. In this regard, another promising direction for future work is the application of our method to classification problems involving image datasets [76, 77, 78, 79, 80, 81, 82, 83, 84], audio datasets [85, 86, 87, 88, 89, 90], and their integration with additional features such as patient information or demographic data [91, 92]. Leveraging datasets that involve more intricate patterns and higher-dimensional data will be instrumental in demonstrating the full potential and scalability of our methodology.
@@ 8 Conclusion
In this paper, we have made several contributions to the fields of binary classification and dimensionality reduction:
  • 1.
    Theoretical insight: We provide a formal framework linking the eigenanalysis of covariance and Hessian matrices to LDA criteria, which enhances the understanding of class separability.
  • 2.
    Novel method: We propose a method that integrates covariance and Hessian matrices for dimensionality reduction, aimed at improving class separability in binary classification tasks.
  • 3.
    Data conditions: We define certain ideal dataset conditions under which our method performs optimally, ensuring that the theoretical assumptions are met and maximizing the method’s effectiveness in practical applications.
  • 4.
    Empirical validation: Our method, tested across various datasets, generally outperforms existing techniques, highlighting its potential robustness and effectiveness under the right conditions.
  • 5.
    Practical utility: The method retains a computational complexity comparable to traditional approaches, suggesting its practicality in real-world applications.
  • 6.
    Improved explainability: By integrating our method with linear SVMs, we enhance the explainability and interpretability of deep neural networks, addressing their typical opacity and facilitating better understanding of their decision-making processes.
@@ 9 Reproducibility statement
The detailed proofs for the theoretical foundations, emphasizing the maximization of between-class mean distance and minimization of within-class variance, are provided in A. The complete source code and datasets for our experiments are accessible through the following links:
@@
  5. 5.
    Illustrative case study on WBCD dataset: Link to Colab notebook
@@
10 Acknowledgments
This research is supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement Sano No 857533 and the project of the Minister of Science and Higher Education ”Support for the activity of Centers of Excellence established in Poland under Horizon 2020” on the basis of the contract number MEiN/2023/DIR/3796.
@@ A.1 Proof of Theorem 1 - Maximizing Covariance for Maximizing Squared Between-Class Mean Distance
For any pair of data points $x_{i}$ and $x_{j}$, where $1\leq i\leq n$ and $n+1\leq j\leq 2n$, the difference between $x_{i}$ and $m_{1}$ is equal to the difference between $x_{j}$ and $m_{2}$, i.e., $x_{i}-m_{1}=x_{j}-m_{2}$. This relationship ensures that the shifted copies of $C_{1}$ and $C_{2}$ maintain the same relative distances from their respective means, preserving the identical distribution in both sets. Furthermore, the variances of $C_{1}$ and $C_{2}$, are equal, denoted as $\sigma_{1}^{2}=\sigma_{2}^{2}=\sigma_{w}^{2}$, and the combined data from $C_{1}$ and $C_{2}$ has a variance of $\sigma^{2}$.
@@
$\sigma^{2}=\frac{1}{2n-1}\left(\sum_{i=1}^{n}(x_{i}-m)^{2}+\sum_{i=n+1}^{2n}(x_{i}-m)^{2}\right)$
@@
$\sigma^{2}$ $=\frac{1}{2n-1}\left(\sum_{i=1}^{n}\left(\left(x_{i}-m_{1}\right)^{2}+\frac{1}{4}d^{2}\right)+\sum_{i=n+1}^{2n}\left(\left(x_{i}-m_{2}\right)^{2}+\frac{1}{4}d^{2}\right)\right)$
@@
$=\frac{1}{2}\left(\sigma_{w}^{2}+\sigma_{w}^{2}\right)+\frac{1}{4}d^{2}$
$=\sigma_{w}^{2}+\frac{1}{4}d^{2}.$
Now, considering the data points representing the projected data onto an (Eigen)vector, we can utilize the Variance Ratio Preservation Theorem for Projection onto a Vector, which establishes the relationship between $\sigma^{2}$ and $\sigma_{w}^{2}$ as follows:
$\sigma_{w}^{2}=\lambda\cdot\sigma^{2}$
@@
$\sigma^{2}$ $=\lambda\cdot\sigma^{2}+\frac{1}{4}d^{2}.$
Let’s rearrange the equation by moving $2\lambda\cdot\sigma^{2}$ to the left side:
$\sigma^{2}-\lambda\cdot\sigma^{2}$ $=\frac{1}{4}d^{2}.$
@@
$(1-\lambda)\cdot\sigma^{2}$ $=\frac{1}{4}d^{2}.$
To solve for $\sigma^{2}$, divide both sides by $(1-\lambda)$:
$\sigma^{2}$ $=\frac{\frac{1}{4}d^{2}}{1-\lambda}=\frac{r^{2}}{1-\lambda}$
@@
We observe that the sign of $\sigma^{2}$ and $d^{2}$ will be the same since the denominator $1-\lambda$ is always positive (as $10<\lambda<$ 1). Therefore, $\sigma^{2}$ is linearly proportional to $d^{2}$.
Hence, maximizing the variance ($\sigma^{2}$) will maximize the squared between-class mean distance ($d^{2}$) as desired.
@@ A.2 Proof of Theorem 2 - Maximizing Hessian for Minimizing Within-Class Variance
However, according to [52], we can approximate the Hessian using Fisher information:
$\mathrm{H}_{\theta}$ $\approx\mathbb{E}_{p_{\theta}}\left[\left(\nabla_{\theta}\log p_{\theta}(c_{i}\mid\theta)\right)^{2}\right]$
@@
where $\mu$ and $\sigma$ are known constants, we can compute the Hessian as follows:
$\mathrm{H}_{\theta}$ $=\mathbb{E}_{p_{\theta}}\left[\left(\nabla_{\theta}\log\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(\theta-\mu)^{2}}{2\sigma^{2}}\right)\right)^{2}\right]$
@@
Combining the previous result with the Hessian calculation, we can conclude that:
$\mathrm{H}_{\theta}$ $=\frac{1}{\sigma^{2}}=\frac{1}{\sigma_{\text{post}}^{2}}$